1st ? - The sum of 2 consecutive integers is -9. Find the numbers.

2nd ? - Tanner has $4.35 in nickels and quarters. If he has 15 more nickels than quareters, how many of each coin does he have?
Thank you for your help!

Answer by stanbon(75887)

1st ? - The sum of 2 consecutive integers is -9. Find the numbers.
1st: x
2nd: x+1
Equation:
x + x+1 = -9
2x = -10
x = -5
x+1 = -4
--------------------------
2nd ? - Tanner has $4.35 in nickels and quarters. If he has 15 more nickels than quareters, how many of each coin does he have?
Quatity Equation: n = q + 15
Value Equation: 5n + 25q = 435 cents
-------------------------------------------
Substitute for "n" and solve for "q":
5(q+15) + 25q = 435
5q + 75 + 25q = 435
30q = 360
q = 12 (# of quarters)
Since n = q + 15, n = 27 (# of nickels)
=========================================
Cheers,
Stan H.